# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### v5

@@ Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of this result. The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2]. The second proof uses improper integrals of formal power series and is due to this author. We explain what “semiformal” means.
1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4]. Hilbert’s proof [5] goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
@@
We may assume that $a_{0}\neq 0$. For $r\in\mathbb{N}$ we define integral polynomials
@@
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof is given in Conrad [3].
Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity. We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ in a way satisfying two conditions.
Only hereditarily at most countable (HMC) sets are directly used in the proof.
No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
We call proofs satisfying these two conditions AB proofs. A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
$x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
the set $x_{n}$ is at most countable (finite or countable). In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
There are at least four ways to an AB proof of the transcendence of $\mathrm{e}$. For the first two see [7]. The other two ways are the topic of the present article. They are two AB proofs of the transcendence of $\mathrm{e}$ based on formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it already in early versions of [7], and we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC. We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$. It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away. We allow such mentions of non-HMC sets in AB proofs for the sake of natural and legible formulations.
In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the uncountable set of maps $f\colon\mathbb{N}_{0}\to R$. They are written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. We work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$. For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
Besides functionals, the set $R[[x]]$ is endowed with the operations of addition and multiplication
${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}}$
which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we form the linear combination
$cf(x)+dg(x)=\sum_{n\geq 0}(ca_{n}+db_{n})x^{n}\,.$}
This is in fact a family of binary operations on $R[[x]]$, one for every pair $c,d$. The structure
R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
is a domain, in which $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
$\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
Less often one uses the (formal) primitive
In this article, $\int$ is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by the convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$. They satisfy for every $c\in K$ that
$\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
In other words, $f(x)\in K\{x\}$ iff $f(x)$ has infinite radius of convergence.
The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient in the product $f\cdot g$,
$[x^{n}](f\cdot g)=\sum_{j=0}^{n}[x^{j}]f\cdot[x^{n-j}]g\,$,}
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations. We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any values of $f(x)$ ($c\in K$)
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}\,.$}
It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
However, we can present two such proofs in the next two sections if formality is relaxed to semiformality. Semiformal functionals are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$. The same holds for semiformal operations. For examples of semiformal functionals and operations on $\mathbb{R}[[x]]$ see Definitions 3.7 and 3.3, respectively. We say that a proof involving elements in $K[[x]]$ is semiformal if it satisfies the following condition.
Only formal and semiformal functionals and operations on $K[[x]]$ are used in the proof. For every formal power series involved in the proof, only at most countably many of its values are used.
In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly present the logical structure of their proof, we reorganize the exposition in [2].
@@
We repeat that the proof in [2] is more general as it establishes the full LW theorem. In that proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is semiformal and not formal because the key player is the infinite sum
$\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal component in an otherwise formal proof.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not really needed in the proof. Condition (B), which states that no formalism of mathematical logic is used, is satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
@@ 3 The proof by semiformal integration
We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [7]. The idea is to translate Hilbert’s proof, described in Section 1, into the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce further functionals and operations on $\mathbb{R}[[x]]$ besides those mentioned in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
@@
We need bounds, obtained in an AB way, on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
@@
The key step in Hilbert’s proof is Euler’s identity. We prove it now in an AB &\& SF way.
@@
We are ready to state Hilbert’s proof in an AB &\& SF way. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an AB &\& SF way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
@@
It remains to bound, in an AB &\& SF way, the numbers
@@
The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and the use of many values of formal power series. Clearly it is an AB proof.

### v6

@@ Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Using formal power series, we then give two algebraic proofs of this result. The first proof is a specialization of the proof of Beukers, Bézivin and Robba of the Lindemann–Weierstrass theorem. The second proof uses improper integrals of formal power series and is due to this author. We explain in what respect both proofs improve upon Hilbert’s proof.
1 Hilbert’s proof and how to improve it
The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4]. Hilbert’s proof [5] goes as follows. We suppose for the contradiction that there exists $n+1$ integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
@@ 1 Hilbert’s proof and how to improve it
We may assume that $a_{0}\neq 0$. For any $r\in\mathbb{N}$ ($\{1,2,\dots\}$) we define integral polynomials
@@
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof appears in Conrad [3].
Hilbert’s proof is a mathematical gem, but there is still a large room for improvement. The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$, which play a key role in the proof, are uncountable sets. Is it possible to give a proof of the transcendence of $\mathrm{e}$ without substantially using uncountable sets? (For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away. Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible, such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]; they actually proved in this way a much more general result, the Lindemann–Weierstrass theorem. What are the main countable players in their proof which replace the functions $x^{n}\mathrm{e}^{-x}$? The complex sequences
$(a_{n})=(a_{0},\,a_{1},\,a_{2},\,\dots),\ a_{n}\in\mathbb{C}\,$,
known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
Why do we hear almost nothing about [2]? (In March 2026 it had $1$ citation in Mathematical Reviews.) In the approach of eliminating uncountable sets from proofs, [2] is a remarkable achievement, even if the authors are probably not aware of it (that approach is not mentioned in [2]). However, only a small minority of professional mathematicians, to which this author belongs, care about the distinction of using or not using uncountable sets in proofs; see, for example, the forthcoming book [7].
The proof of Beukers, Bézivin and Robba deserves a new edition, so to say, and we review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$. In Section 3 we present another uncountable-sets-free proof of the transcendence of $\mathrm{e}$ using formal power series. It was found by this author and is based on the natural idea of translating Hilbert’s proof to the language of formal power series.
To conclude the Introduction, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
whenever $b_{j}$ and $\alpha_{j}$ are $2t\geq 2$ algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly present the logical structure of the proof, we reorganize the exposition in [2].
@@
In the original proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
@@ 3 The proof by semiformal integration
We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence. For $f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
@@
We call these operations on $\mathbb{R}[[x]]$ semiformal because coefficients in them arise by limit transitions; in the standard formal operations they are obtained by finite expressions. The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
@@
We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
@@
We prove Euler’s identity without using uncountable sets.
@@
We are ready to state Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an uncountable-sets-free way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we compute semiformally.
@@
It remains to bound, in an uncountable-sets-free way, the numbers